OpenStudy (girlgoyleh.):

What is the constant of variation for the quadratic variation? 6y = 4x^2 A.3/2 B.4/9 C.2/3 D.-2

OpenStudy (anonymous):

solve for y divide by 6 for both sides y=4/6x^2 y=2/3x^2

OpenStudy (anonymous):

constant variation is the slope which 2/3

OpenStudy (girlgoyleh.):

thank you, i just wanted to make sure. thanks for explaining it :)